OpenStudy (konradzuse):

integral of x/((3-x^2) ^2)

OpenStudy (lgbasallote):

let u= \(3 - x^2 \rightarrow du = -2xdx\)

OpenStudy (lgbasallote):

therefore you have \[\frac 12 \int \frac{du}{u^2}\]

OpenStudy (konradzuse):

I got \[\int\limits (x) * (3-x^2) ^{-1/2} dx\] u = 3-x^2 du = 2x dx 1/2du = xdx \[1/2 \int\limits (u)^{-1/2} + c\] \[(1/2) * (-2/3) u^{-2/3}\] Which should equal\[-{2/6} (3 - x^2)^ {-3/2}\]

OpenStudy (lgbasallote):

waut...why -1/2??

OpenStudy (lgbasallote):

isnt that squared only?

OpenStudy (konradzuse):

At first I thought it wasn't negative, but then I thought it does that when it goes from the denom to the num?

OpenStudy (lgbasallote):

i meant why 1/2

OpenStudy (lgbasallote):

the denominator is squared right?

OpenStudy (konradzuse):

I'm sorry it was supposed to be \[\sqrt{3-x^2}\]

OpenStudy (konradzuse):

I messed up :(

OpenStudy (lgbasallote):

your process is all right except the final answer

OpenStudy (lgbasallote):

1) simplify 2/6 into 1/3 2) where did 3/2 come from?

OpenStudy (lgbasallote):

oh you miswrote the previous step as 2/3 instead of 3/2

OpenStudy (lgbasallote):

nevermind 3/2 is correct

OpenStudy (konradzuse):

Yeah I messed up, I had different stuff written at first and messed up everything else sorry.

OpenStudy (lgbasallote):

in simplified term it should look like \[-\frac{1}{3(3-x^2)^{3/2}}\]

OpenStudy (konradzuse):

now that I think about it shouldn't it be -1/2 + 2/2 to be Positive 1/2?

OpenStudy (lgbasallote):

good point

OpenStudy (lgbasallote):

how could i have overlooked that lol

OpenStudy (lgbasallote):

so 1/2 cancels leaving you with \((3-x^2)^{1/2}\) only

OpenStudy (konradzuse):

so it should be \[-(1/3) * (3 - x^2)^{1/2} + C\]

OpenStudy (lgbasallote):

-1/3? no

OpenStudy (lgbasallote):

integral of (3-x^2)^-1/2 is \(2(3 - x^2)^{1/2}\) agree?

OpenStudy (konradzuse):

oh yeah woopsies, keep confusing myself lol. I just looked back at your last post.

OpenStudy (lgbasallote):

so then there's a 1/2 in the outside..2 just cancels

OpenStudy (konradzuse):

\[(3−x^2)^{1/2}+C\]

OpenStudy (lgbasallote):

^that

OpenStudy (konradzuse):

apparently that isn't correct, I'm going to try this over from scratch.

OpenStudy (lgbasallote):

no?!

OpenStudy (lgbasallote):

lemme clarify this...this is the question? \[\int \frac{x}{\sqrt{3-x^2}}\]

OpenStudy (konradzuse):

yup

OpenStudy (lgbasallote):

OHHH of course *facepalm* the negative sign!!

OpenStudy (lgbasallote):

it's \[-\frac{1}{2} [2(3-x^2)^{1/2}]\] there should be a negative sign

OpenStudy (lgbasallote):

sorry for overlooking a lot of things :S

OpenStudy (konradzuse):

oopsies :P No it's okay, I appreciate all of your help :).

OpenStudy (konradzuse):

I have messed up badly on this question :P so \[-(3-x^2) ^{1/2}\]

OpenStudy (lgbasallote):

+C :P lol

OpenStudy (konradzuse):

That's an obvious :P. By the way when do we actually find what C is? I feel like it's always something we write, but never figure out...

OpenStudy (konradzuse):

Now that I think about it how did we get -1/2? Was it because 3-x^2 was -2x dx?

OpenStudy (konradzuse):

That would make sense I think I overlooked it :)

OpenStudy (lgbasallote):

yes derivative of 3-x^2 is -2xdx

OpenStudy (lgbasallote):

and you will know what C is if there are numbers on the integral like \[\int_2^1 x^2dx\] don't mind that for now lol

OpenStudy (konradzuse):

Ah okay, I've dealt with those already, I figured there was somehing else to it, thanks ! :).